The perimeter of an isosceles triangle is 37 inches. the lenghts of the two equal legs are 6 inches less than 3 times the length of the base. find the lengths of the three sides.
Answer by checkley75(3666)

SIDES=3B-6
2(3B-6)+B=37
6B-12+B=37
7B=37+12
7B=49
B=49/7
B=7 in. FOR THE BASE LENGTH,
SIDE=3*7-6
SIDE=21-6
SIDE=15 in.
PROOF
2*15+7=37
30+7=37
37=37